A radiator contains 14 quarts of fluid, 25% of which is antifreeze. How much fluid should be drained and replaced with pure antifreeze so that the new mixture is 50% antifreeze? (Round your answer to two decimal places.)

Answer by mananth(16946)

---------percent ---------------- quantity
Antifreeze 100 ---------------- x quarts
Fluid 25---------------- 14 - x quarts
Mixture 50 ---------------- 14

100 x + 25 ( 14 - x ) = 50 * 14

100 x + 350 - 25 x = 700
100 x - 25 x = 700 - 350
75 x = 350
/ 75
x=4.67 quarts 100.00% Antifreeze
9.33 quarts 25.00% Fluid
4.67 quarts has to be drained